what are the steps to solving this equation:
2v-3 over 5 is greater than or equal to 7

Answer by Alan3354(69443)

(2v-3)/5 >= 7
1st, it's not an equation, it's an inequality. The 2 sides are not equal.
(2v-3)/5 >= 7
Multiply both sides by 5
2v - 3 >= 35
Add 3
2v >= 38
Divide by 2
v >= 19
---------
Was that hard?